Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows: Sales $ 2,080,000 Variable expenses 1,040,000 Contribution margin 1,040,000 Fixed expenses 180,000 Net operating income $ 860,000

5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $69,000 increase in advertising, would increase this year's unit sales by 25%. a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented? b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?
Answer #1

Particulars

per unit

Units

Amount ($)

A

Sales

69.6

32500

2,262,000

B

variable expenses

40

32500

1,300,000

C=A-B

Contribution margin

29.6

32500

9,62,000

D

Fixed expenses

180,000

E

Advertising

69,000

F=C-D-E

Net Operating Income

$ 713,000

Number of units sold = 2080,000/80 = 26,000 units

Adjusted number of units sold = 26000 + 25% of 26,000 = $ 32,500

New selling price = 80 – 13% of 80 = $ 69.6

Advertising cost is added expenditure of amount $ 69,000

  1. If his ideas are implemented, the firm would earn net operating income of $ 713,000
  2. Net operating income would decrease by 860,000 – 713,000 = $ 147,000.